OpenStudy (anonymous):

Belle's Bake Shop makes croissants that cost $1.75 each. Past experience shows that 10 percent of croissants will spoil and have to be discarded. Assuming Belle wants a 45 percent markup based on cost and produces 300 croissants, each croissant should sell for: (Points : 4)

OpenStudy (anonymous):

OK, so of the 300 croissants, 10% will spoil. That is, only 90% of the 300 will be sell-able. Numerically this can be described as 90%*300 which is 270.

OpenStudy (anonymous):

45% markup means the same thing as 145% of the original price, yes? So, 270 croissants at 145% of 1.75 can be described numerically like this: 275*145%*1.75 which is the same as 275*1.45*1.75 The above will get you the total monetary value of selling all 270 croissant. However, we want only how much each croissant will cost. So we divide 270*1.45*1.75 by 270, and the new expression is 1.45*1.75.

OpenStudy (anonymous):

I don't think something's right here, but I'm not sure what. It feels like we need to incorporate my first post....tell me what you think

OpenStudy (pfenn1):

You have to take into account that you made 300 croissants at a cost of $1.75 each but you are only able to sell 90% of them or 270. So you want to sell each croissant at a price where 270 x price =1.45 x cost to make 300 bagels or:\[270 \times P=1.45(300 \times 1.75)\]